OpenStudy (anonymous):

Evaluate: 48 ÷ (6 + 2) · 3

OpenStudy (phi):

what is the first thing you do first?

OpenStudy (anonymous):

We use BODMAS ( Brackets, Of , Division, Multiplication, Addition,Subtraction) to decide the preference. So 48÷(6+2)• 3 =48÷(8)• 3 =6• 3 =18

OpenStudy (anonymous):

I did that but for some reason it came out wrong. I did PEMDAS. perenthesis, exponents, Multiplication, Division, & subtraction

OpenStudy (anonymous):

you use bodmas

OpenStudy (anonymous):

its the Same thing :p

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you though ☺

OpenStudy (anonymous):

u welcom

OpenStudy (phi):

I did that but for some reason it came out wrong. I did PEMDAS. Order also matters. You do these left to right. After doing the sum inside the parens you get 48 ÷ 8 · 3 now do it left to right: (48 ÷ 8) · 3

OpenStudy (anonymous):

right. But wouldnt the numbers in the brackets go first ?